OpenStudy (rath111):

Prove that the given equations are identities: 1+tan(theta)tan2(theta)=tan2(theta)cot(theta)-1

OpenStudy (lgbasallote):

do you know the equation for \(\tan 2\theta\)? coz i have forgotten and i might be able to help if i know it

OpenStudy (rath111):

Yes I do know. Hold on...

OpenStudy (rath111):

tan2(theta)=2tan(theta)/1-tan^2(theta)

OpenStudy (lgbasallote):

gimme a minute to figure this out...

OpenStudy (rath111):

Okay.

OpenStudy (lgbasallote):

have you tried it? if so how far did you get?

OpenStudy (rath111):

I've tried it but... I got to 1+tan(theta)tan2(theta)=tan^2(theta)/1-tan^2(theta) Sure I did something wrong though...

OpenStudy (lgbasallote):

hmm you did the right hand side huh....i got sec^2 theta/1 - tan^2 theta for that

OpenStudy (rath111):

Someone was typing a reply, then he went silent.

OpenStudy (lgbasallote):

\(\LARGE \frac {2 \tan \theta \cot \theta}{1 - \tan^2 \theta} - 1\) \(\LARGE \frac {2 (\frac{\cancel{\sin \theta}}{\cancel{\cos \theta}}) (\frac{\cancel{\cos \theta}}{\cancel{\sin \theta}})}{1 - \tan^2 \theta} - 1\) \(\LARGE \frac {2 }{1 - \tan^2 \theta} - 1\) \(\LARGE \frac {2 }{1 - \tan^2 \theta} - 1\) lol this is what i got...ignore the sec something i said a while ago

OpenStudy (rath111):

I don't know... What? I'm confused.

OpenStudy (anonymous):

1+tan(theta)tan2(theta)=tan2(theta)cot(theta)-1 ok so lets do right side first and change everything to sin and cos like we did before okay so tan2(theta)=2tan(theta)/1-tan^2(theta) and cot theta = costheta/sintheta ok so i am going to change theta into x's alright tan2xcotx - 1 = so this is just (sin2x/cos2x)(cosx/sinx) - 1 = (2sinxcosx/cos^2x - sin^2x)(cosx/sinx) - 1 = (2cos^2x/cos^2x - sin^2x) - 1 = (2cos^2x - cos^2x +sin^2x)/cos^2x - sin^2x the top just becomes a 1 so now we got 1/cos2x for the right side okay :)

OpenStudy (anonymous):

now for the left side 1+tan(theta)tan2(theta) change everything to sin and cos alright so we got 1+ (tanx)(tan2x) and that is 1+ (sinx/cosx)[(2sinxcosx/(cos^2x - sin^2x)] = 1 + [2sin^x/(cos^2x - sin^2x)] = [(cos^2x - sin^2x + 2sin^2x)/(cos^2x - sin^2x)] the top just becomes a 1 so you got 1/cos2x just like we wanted :) alright @Rath111

OpenStudy (paxpolaris):

\[RHS=\LARGE \frac {2 }{1 - \tan^2 \theta} - 1={2-1+\tan^2(\theta) \over 1-\tan^2(\theta)}\] then do the same with LHS

OpenStudy (lgbasallote):

@PaxPolaris correct me if im wrong but in proving i think you are only allowed to touch one side

OpenStudy (paxpolaris):

no you can simplify both sides, if you need to.

OpenStudy (paxpolaris):

you could also first re-write the original equation into: \[\Large 2= \tan(2\theta)\left[ \cot(\theta)- \tan(\theta)\right]\]

OpenStudy (rath111):

Trying to absorb this information...

OpenStudy (rath111):

The left side is what's so trippy right now...

OpenStudy (rath111):

k getting it...

OpenStudy (rath111):

That took forever for my weak brain to process, but I finally understood it after messing with the problem and looking back and forth with what I thought was the best answer. Thank you :)